OpenStudy (crashonce):

The four positive numbers 3, x, y, 9 are such that 3, x, y form a GP and x, y, 9 form an AP. FInd x+y. @ganeshie8

OpenStudy (crashonce):

@perl

OpenStudy (campbell_st):

find the common ratio for the 3 term GP\[r = \frac{x}{3} = \frac{y}{x}\] cross multiply and you get equation 1 \[\frac{x^2}{3} = y\] now look at the 3 term AP and find the common difference \[y - x = 9 - y~~~ or ~~~~~ 2y - x - 9 = 0\] call this equation 2 now you need to solve simultaneous equations using the substitution method substitute equation 1 into equation 2 \[2(\frac{x^2}{3}) - x - 9 = 0\] multiply every term by 3 \[2x^2 - 3x - 27 = 0\] now you have an equation that can be solved by factoring eliminate the negative zero, and only consider the positive when you get x, substitute into equation 1 or 2 and solve for y. hope it helps

OpenStudy (anonymous):

3,x,y form a GP => x^2 = 3y x,y,9 form an AP => 2y=9+x => x^2 = 3(9+x)/2 => 2x^2 -3x -27 = 0 => 2x^2 + 6x -9x -27=0 => (2x-9)(x+3) = 0 x=9/2 or x =-3 2y = 9 + 9/2 => 4y = 27 y = 27/4 x+y = 9/2 + 27/4 = 45/4 2y = 6 => y = 3 x+y = -3 + 3 =0 Hope its Correct...

OpenStudy (crashonce):

0 isnt an answer option the options are 9.5, 10, 10.5, 11.25, 13.5 @campbell_st could you check yourss cos i dont think its right

OpenStudy (anonymous):

45/4 is 11.25 which is the answer @CrashOnce :)

OpenStudy (anonymous):

Answer set is 11.25 or 0

OpenStudy (crashonce):

r u sure?, i cant seem to prove that is correct...

OpenStudy (anonymous):

3, 9/2, 27/4, 9 is the series... => (9/2)^2 = 3*27/4 (GP) and => 2*27/4 = 27/2 and 9/2 + 9 = 27/2 (AP) So its Correct...

OpenStudy (campbell_st):

ok... so what positive solution did you get for x...?

OpenStudy (campbell_st):

@sriramkumar the aim of open study is to assist learning and not to give answers.

OpenStudy (crashonce):

ok thanks both of u, though @campbell_st sriamkumar provided working which was helpful

OpenStudy (campbell_st):

lol.... well I got you to the quadratic... if you are working at this level I would have expected you co factor the quadratic.... as I said before, the aim is to guide you... not to answer homework or exam questions for you.

OpenStudy (crashonce):

sorry bout that but thanks both of u